OpenStudy (anonymous):

Another Algebra Problem! -6(9x - 6) = 306 Help plz!

OpenStudy (anonymous):

I think i got it. I came up with -5!

OpenStudy (anonymous):

-17

OpenStudy (anonymous):

-54x+36=306 -54x=270 x=306/(-54) x=-5

OpenStudy (anonymous):

You didnt distribute .

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Solve for x: -6 (9 x-6) = 306 Factor the polynomial, -6+9 x. Factor 3 from the polynomial 9 x-6: -63 (3 x-2) = 306 Multiply -6 and 3 together. -6×3 = -18: -18 (3 x-2) = 306 Divide both sides by a constant to simplify the equation. Divide both sides of -18 (3 x-2) = 306 by -18: (-18 (3 x-2))/(-18) = 306/(-18) Any nonzero number divided by itself is one. (-18)/(-18) = 1: 3 x-2 = 306/(-18) Reduce 306/(-18) to lowest terms. The gcd of 306 and -18 is 18, so 306/(-18) = (18×17)/(18 (-1)) = 18/18×17/(-1) = 17/(-1): 3 x-2 = 17/(-1) Simplify the sign of 17/(-1). Multiply numerator and denominator of 17/(-1) by -1: 3 x-2 = -17 Isolate terms with x to the left hand side. Add 2 to both sides: 3 x+(-2+2) = -17+2 Look for two terms that sum to zero. -2+2 = 0: 3 x = -17+2 Evaluate -17+2. -17+2 = -15: 3 x = -15 Solve for x. Divide both sides of 3 x = -15 by 3: (3 x)/(3) = (-15)/3 Any nonzero number divided by itself is one. 3/3 = 1: x = (-15)/3 Reduce (-15)/3 to lowest terms. The gcd of -15 and 3 is 3, so (-15)/3 = (3 (-5))/(3×1) = 3/3×-5 = -5: Answer: | | x = -5

OpenStudy (anonymous):

Hopefully that helps.

OpenStudy (anonymous):

A lot Ty :D

OpenStudy (anonymous):

@BabyPhoenix just stick with @rajatgang answer